OpenStudy (anonymous):

Solve for n: 7n - (4 - 2n) = 3(n + 5) - 1

OpenStudy (anonymous):

um i would like to buy a vowel

OpenStudy (anonymous):

All you have to do is distribute the negative and the 3

OpenStudy (anonymous):

Sure what vowel

OpenStudy (anonymous):

always remember to combine like terms

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

\[7n-(4-2n)=3(n+5)-1\]\[7n-4+2n=3n+15-1\]\[9n-4=3n+14\]\[6n=18\]\[n=3\] To check: \[7(3)-(4-2(3))=3((3)+5)-1\]\[21-(4-6)=3(8)-1\]\[21-(-2)=24-1\]\[23=23\]CHECK